OpenStudy (anonymous):

d-1 over 6(fraction) = d-3 over 3(fraction)

OpenStudy (anonymous):

\[(d-1)\div6 = (d-3)\div3\]

OpenStudy (anonymous):

Any ideas where to start?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

Ok, what is your idea?

OpenStudy (anonymous):

mulitiple the first fraction by its denominator and the second fraction by the denominator

OpenStudy (anonymous):

great, we can do that, what do we get if we multiply both sides by 6?

OpenStudy (anonymous):

I will be right back.

OpenStudy (anonymous):

d-1=d-3/3*6